Write the equation in standard form using only integers.
y = 2/3 x -4

Answer by ramanna(5)

Given equation is
y=2/3x-4
first we bring it to the standard form by taking x to the left of equal sign,
this becomes now as,
-2/3x+y=-4
now the coefficient of x is not an integer, we need to make it an integer, for that multiply the whole equation by 3, then the equation becomes..
-2x+3y=-4
this equation is now in standard form.